### v2

# VerifyBench: A Systematic Benchmark for Evaluating Reasoning Verifiers Across Domains

Abstract
Large language models (LLMs) increasingly rely on reinforcement learning (RL) to enhance their reasoning capabilities through feedback. A critical challenge is verifying the consistency of model-generated responses and reference answers, since these responses are often lengthy, diverse, and nuanced. Rule-based verifiers struggle with complexity, prompting the use of model-based verifiers. However, specialized verifiers lack flexibility, while general LLM judges can be inconsistent. Existing research primarily focuses on building better verifiers, yet a systematic evaluation of different types of verifiers’ performance across domains remains lacking, severely constraining the reliable development of Reinforcement Learning with Verifiable Reward (RLVR). To address this, we propose VerifyBench–a cross-domain comprehensive benchmark for systematically evaluating verifiers. We construct 4,000 expert-level questions covering mathematics, physics, chemistry, and biology. Each question is equipped with reference answers and diverse responses. The reliability of the evaluation is ensured through a rigorous annotation process conducted by a multidisciplinary expert team. We design a four-dimensional experimental framework to comprehensively compare the performance boundaries of specialized verifiers and general LLMs under combined conditions of extracted answers vs. complete responses, and short vs. long outputs. Our evaluation uncovers fundamental trade-offs in verifiers: while specialized verifiers achieve leading accuracy (the best model reaching 96.48% in chemistry), they exhibit deficiencies in recall; general models show stronger inclusivity but unstable precision. More importantly, we discover verifiers’ high sensitivity to input structure and inherent limitations in cross-domain generalization, providing critical insights into the bottlenecks of current verifier technology.
Introduction
Large language models (LLMs) (Guo et al. 2025; Achiam et al. 2023; Team et al. 2025) have achieved significant breakthroughs in complex reasoning, planning, and symbolic problem-solving (Cao et al. 2025). This progress is largely driven by Reinforcement Learning with Verifiable Reward (RLVR) (Luo et al. 2025; Zeng et al. 2025), which employs an external feedback loop to refine policy model. As depicted in Figure 1, this mechanism enables LLMs to generalize and produce high-quality responses. In recent work, this paradigm has been extended beyond rule-based reward (Hynek Kydlíček 2024) to include verifier-based learning (Chen et al. 2025; Ma et al. 2025a), where model outputs are evaluated by a separate verifier model.
However, the inherent unreliability of current verifier systems remains a critical challenge (Huang et al. 2025). Rule-based verifiers, relying on rigid pattern matching, are limited in generalization and brittle with diverse, nuanced LLM responses. They frequently misclassify semantically correct but non-canonical responses, especially in complex domains like mathematics and science. This fundamental limitation creates misalignment between model performance and reward signals, directly hindering RLVR’s scalability and trustworthiness in real-world deployment.
To overcome these constraints, model-based verifiers have emerged, including specialized models (finetuned on labeled data) and general-purpose LLMs (Yang et al. 2025; Cai et al. 2024; Qwen et al. 2025) acting as judges. While promising enhanced flexibility, their practical application introduces new issues. Specialized verifiers, though accurate, often lack adaptability to novel expressions. General-purpose LLMs frequently suffer from inconsistency and insufficient precision, particularly for step-by-step reasoning (Huang et al. 2025). Crucially, despite intense focus on verifier development, a systematic and comprehensive evaluation of diverse verifier types across varied domains and conditions remains absent. This significant gap impedes robust RLVR development, leaving practitioners without clear guidance.
To fill this critical void, we introduce VerifyBench: a cross-domain comprehensive benchmark for systematic evaluation of verifiers. VerifyBench is constructed from 4,000 expert-level questions spanning mathematics, physics, chemistry, and biology. Each question includes reliable reference answers and diverse Chain-of-Thought (CoT) (Ma et al. 2025b; Wei et al. 2022) responses. Gold-standard judgment labels are established through a rigorous, fine-grained human annotation process by a multidisciplinary expert team, ensuring unparalleled reliability. This enables systematic analysis of verifier behavior across a four-dimensional experimental framework, varying input granularity (boxed final answers vs. full reasoning traces) and output constraints.
Our benchmark and experimental design are driven by three primary goals. First, to provide a systematic, multidisciplinary evaluation platform for verifiers across STEM domains. Second, to comprehensively analyze intricate behavioral differences between verifier types. Third, by simulating realistic RLVR deployment scenarios through varied input and output conditions, we uncover how these factors affect verifier reliability and expose critical bottlenecks in their current design.
Through controlled experiments, we rigorously evaluate both specialized verifiers (finetuned LLMs) and general-purpose LLMs (zero-shot or few-shot verifiers). Our findings reveal fundamental trade-offs in verifier design: specialized models offer leading precision but exhibit deficiencies in recall and struggle with diverse expressions. Conversely, general LLMs with larger model sizes demonstrate high inclusiveness but suffer from inconsistent structured judgment and a heightened risk of false positives. Importantly, we empirically uncover verifiers’ acute sensitivity to input structure and inherent limitations in cross-domain generalization, providing critical insights into the fundamental challenges facing current verifiers. VerifyBench offers a rigorous foundation for developing trustworthy evaluations in RL-trained LLMs.
Our main contributions are as follows:
We introduce VerifyBench, a novel, multidisciplinary benchmark of 4,000 expert-level questions across mathematics, physics, chemistry, and biology, featuring fine-grained human annotations and diverse CoT responses from state-of-the-art LLM, specifically designed for systematic verifier evaluation.
We conduct a comprehensive empirical study comparing various verifier types—specialized models vs. general-purpose LLMs—under varying input contexts and output constraints. This systematic framework reveals fundamental judgmental trade-offs and highlights their limitations, particularly concerning precision, recall, and robustness to diverse reasoning.
We identify key challenges critical for verifier deployment, including the pervasive accuracy-recall trade-offs, input structure sensitivity, and cross-domain generalization limitations. Based on these insights, we propose actionable directions for designing more robust and generalizable verification systems to accelerate the reliable development of RLVR.
Related Work
Reinforcement Learning with Verifiable Reward
Reinforcement learning (RL) (Hu et al. 2025; Yu et al. 2025; He et al. 2025; Yue et al. 2025) has emerged as a powerful paradigm for aligning LLMs with task-specific goals by optimizing reward signals through interaction. In tasks with well-defined outputs—such as mathematics (Hendrycks et al. 2021), programming (Jain et al. 2024), and logic puzzles (Xie et al. 2025)—automatic verifiers are frequently employed to provide reward feedback by evaluating the correctness of model responses. This technique enables scalable supervision without exhaustive human labeling and has been integrated into the training pipelines of many recent high-performing models (Li, Zou, and Liu 2025b). While effective in principle, the choice and reliability of the verifier are critical. Most existing systems rely on rule-based verifiers that operate by matching the model’s final answer against a reference (Li, Zou, and Liu 2025a), often using hand-crafted rules or symbolic equivalence criteria. These approaches are efficient but brittle: they may fail to recognize semantically correct answers expressed in alternative forms or with minor formatting differences. Furthermore, in multi-step reasoning scenarios, a strict match on the final boxed answer can overlook the model’s overall process quality. This creates a mismatch between verifier judgment and human evaluation standards, especially when the reasoning path is correct but the expression is unconventional.
Model-based Verifier
To address the limitations of static rule-based evaluation (Hynek Kydlíček 2024), recent work explores trained verifier (Chen et al. 2025; Huang et al. 2025) or general LLMs (Qwen et al. 2025; Yang et al. 2024) as verifiers. Specialized verifiers are trained to predict whether a model-generated answer is valid based on responses and reference answers. These verifiers can handle richer linguistic and symbolic variation, and offer more flexible evaluation in tasks involving complex reasoning (Jiang et al. 2025). However, trained verifiers may be overly tolerant, leading to reduced selectivity and accuracy (Ma et al. 2025a). Some studies also propose using general LLMs as verifiers (Jin et al. 2025), leveraging their generalization capabilities and world knowledge in a few-shot or zero-shot setting. Despite their increasing use, a systematic understanding of how different types of verifiers perform under varying input and output conditions remains lacking. Our work addresses this gap through a unified benchmark and controlled comparison across verifier paradigms.
Crucially, while prior research has focused on developing specific verifier models, a comprehensive and systematic evaluation of their performance across diverse scenarios and domains is absent. Our work, VerifyBench, uniquely addresses this gap by establishing a rigorous benchmark and framework to assess existing and future verifiers, rather than proposing new architectures.
Task Overview
[…]
denotes the -th question.
denotes the model’s complete response to question.
denotes the reference answer (i.e., ground truth).
For each instance , the verification process is formally defined by a verifier function , which produces a binary evaluation result :
[…]
is the binary judgment function. It takes the processed model output from and the reference answer as input, returning if the model’s response is deemed consistent with the reference answer, and otherwise.
VerifyBench
Question Collection and Solution Generation
We curate a total of 4,000 high-quality, expert-level questions across four disciplines: mathematics, physics, chemistry, and biology (1,000 per subject). These questions are designed to be challenging, incorporating rich domain-specific background, requiring advanced reasoning capabilities from LLMs, and allowing for open-ended, expressive answer spaces—thereby increasing benchmark complexity. Our definition of “expert-level” questions refers to problems typically encountered in university-level coursework or national/international academic competitions, characterized by their requirement for multi-step reasoning and synthesis of multiple concepts.
As shown in Figure 2, we establish a dedicated data team composed of multidisciplinary experts with strong academic backgrounds in each respective domain to ensure data quality. The collection process strictly adheres to rigorous quality control standards and follows several key steps:
Drawing from both domestic and international sources, team members manually select problems that align with our “expert-level” definition. These questions are chosen to ensure broad topical coverage, high cognitive demand, and diversity in conceptual, theoretical, and applied reasoning, always paired with correct and clear reference answers. A detailed description of our question selection criteria and source identification process is provided in Appendix.
Each question undergoes 2 rounds of manual review. Cross-validation is conducted by graduate-level (Master’s and PhD) reviewers to identify redundant knowledge points, overly simplistic phrasing, ambiguous wording, or irrelevant content. This careful review process ensures that the questions and reference answers are professional, coherent, and free from noise. Further specifics on our review protocols and reviewer qualifications can be found in Appendix.
[…]
Benchmark Annotation
To ensure high-quality and consistent labeling of alignment between model responses and reference answers, we adopt a two-stage annotation process.
In the first stage, each instance (question, complete model response, and reference answer) is annotated by two annotators and reviewed by an independent reviewer. In the second stage, 200 questions are randomly sampled from each discipline (800 in total across 4 domains) for cross-validation by two annotators. Throughout the annotation process, we implement a real-time feedback mechanism: for complex, ambiguous, or disputed cases, annotators consult the data construction and standards team, reach a consensus through discussion, and document representative cases to iteratively refine the annotation guidelines.
For short-answer evaluation, annotation criteria include:
1.
Recognition of superficial answer variations, such as case insensitivity (“a” vs. “A”) or equivalence of “alpha” and “”.
2.
Equivalence judgments across LaTeX expressions, symbolic formats, and natural language descriptions.
3.
[…]
4.
Integrative judgment combining the reference answer’s explanatory context, question requirements, and the model’s response.
5.
Cases where the model derives the correct result but outputs an incorrect final answer are marked as wrong.
Given the domain-specific nuances, we further extend and refine annotation principles as follows:
Mathematics and Physics:
Attention is paid to units, notation, and equivalent solution paths. In physics, dimensional consistency and valid conversions are emphasized.
Chemistry:
Chemical names (e.g., NaCl and sodium chloride), formulas, and states are treated as equivalent. Both International Union of Pure and Applied Chemistry (IUPAC) (IUPAC 1992) and common names are accepted, with flexibility in reaction equation formats.
Biology:
Variations in terminology (technical, vernacular, Latin) are tolerated if biologically accurate. Emphasis is placed on conceptual validity and mechanistic correctness.
Technical Formatting and Language:
Tolerances are defined for spelling, punctuation, and formatting. In long-chain reasoning, correctness of logic and final answer are prioritized over surface-level language differences.
Benchmark Statistics
This section offers a detailed summary of VerifyBench, highlighting its scale, composition, and core characteristics across disciplines. The benchmark contains 4,000 questions, each paired with a reference answer and a model-generated response, intended to support comprehensive and rigorous verifier evaluation. To ensure balanced representation, we include 1,000 questions from each of four major scientific domains: mathematics, physics, chemistry, and biology. Key statistics are summarized in Table 1.
Statistic | Value |
Total Questions | 4,000 |
Questions per Discipline | 1,000 |
Average Question Length | 186 tokens |
Average Model Response Length | 4,553 tokens |
Total Annotated Instances | 4,000 |
Label Distribution (Correct / Incorrect) | 45% / 55% |
Inter-Annotator Agreement (IAA) | 0.88 – 0.92 |
Verifier Math Chemistry Biology Physics Overall Trained Verifier xVerify-0.5B-I 79.60%\64.92% 94.77%\96.50% 86.75%\92.86% 92.48%\94.86% 88.77%\88.66% xVerify-3B-Ia 81.28%\56.91% 94.77%\96.21% 89.56%\90.18% 91.57%\92.24% 89.24%\85.35% xVerify-8B-I 81.58%\55.80% 96.39%\96.50% 87.95%\91.07% 92.68%\92.69% 90.03%\85.47% xVerify-9B-C 83.03%\54.14% 96.28%\94.46% 89.96%\91.07% 93.78%\92.81% 90.96%\84.76% general-verify - - - - - R1-Distill-Verifier-1.5B - - - - - General LLM as Judge Qwen2.5-7B-Instruct 78.18%\91.99% 89.45%\82.22% 68.67%\83.04% 86.66%\98.17% 83.52%\93.74% Qwen2.5-14B-Instruct 78.40%\90.61% 90.05%\82.80% 54.62%\83.04% 91.28%\96.46% 84.10%\91.55% Qwen2.5-32B-Instruct 80.98%\76.24% 85.63%\61.81% 61.45%\66.96% 95.39%\97.26% 85.34%\83.58% Qwen3-8B/14B/32B - - - - - DS-R1-Distill-Qwen-7B/14B/32B - - - - -
VerifyBench emphasizes long-form, multi-step reasoning, representing a substantial challenge for existing verifiers. The high average token count for both responses (Table 1) reflects the benchmark’s focus on detailed, explanatory solutions and verbose model outputs. Our two-stage human annotation pipeline ensures high-quality binary labels, with inter-annotator agreement (IAA) (Artstein 2017) scores ranging from 0.88 to 0.92, confirming annotation reliability.
Importantly, VerifyBench is designed to expose verifier limitations in both accuracy and robustness. It includes:
Semantically diverse correct responses, which require verifiers to go beyond string-level matching;
Subtle reasoning errors and logical inconsistencies, testing fine-grained discrimination;
Cross-domain coverage, requiring awareness of domain-specific notation, units, and reasoning conventions.
Together, these characteristics create a rigorous and realistic evaluation setting that challenges verifiers to combine natural language understanding with domain-specific reasoning competence. We provide a physics case in Figure 4. Appendix provides more detailed cases.
Experiment
Settings
Response Generation.
To simulate realistic response patterns of reasoning-capable language models, we employ QwQ-32B (Team 2025) to generate detailed CoT responses for all 4,000 multi-subject questions in VerifyBench. We use the official recommended hyperparameters: temperature 0.6, TopP 0.95, MinP 0, TopK 40, and no repetition penalty. The maximum response length for QwQ-32B is set to 32,768 tokens.
Baseline.
We evaluate two broad categories of LLMs as verifiers. The first category comprises general-purpose open-source LLMs, including the Qwen2.5-Instruct series (7B, 14B, 32B) (Qwen et al. 2025), the Qwen3 series (8B, 14B, 32B) (Yang et al. 2025), and the DeepSeek-R1-Distill-Qwen series (7B, 14B, 32B) (Guo et al. 2025). The second category consists of specialized verifiers that have been trained on large-scale verification datasets, such as the xVerify series (0.5B, 3B, 8B, 9B) (Chen et al. 2025), general_verifier(Ma et al. 2025a), and R1-Distill-Verifier-1.5B (Huang et al. 2025).
Among these, models from the xVerify and Qwen2.5-Instruct series directly output binary verification decisions (e.g., “correct” or “incorrect”). In contrast, generative models—such as the Qwen3 and DeepSeek-R1-Distill-Qwen series, as well as general_verifier and R1-Distill-Verifier-1.5B—produce reasoning traces alongside their final judgments.
Verifier Inference.
The prompts for all models are unified as the prompt of xVerify (Chen et al. 2025). All models are evaluated under the same decoding configuration: temperature is set to 0.2, and TopP to 0.95.
Evaluation Variants.
We consider the following four experimental settings:
[…]
Main Results
Verifier Math Chemistry Biology Physics Overall Trained Verifier xVerify-0.5B-I 79.38%\54.14% 94.67%\94.46% 87.15%\91.07% 92.38%\92.81% 88.64%\84.76% xVerify-3B-Ia 81.18%\56.91% 95.03%\96.21% 88.76%\90.18% 91.88%\92.24% 89.38%\85.35% xVerify-8B-I 81.58%\55.80% 96.28%\96.50% 89.16%\91.07% 92.58%\92.69% 90.06%\85.47% xVerify-9B-C 82.78%\64.92% 96.18%\96.50% 89.96%\92.86% 93.98%\94.86% 90.86%\88.66% general-verify 68.77%\88.12% 75.13%\88.63% 73.09%\85.71% 94.32%\97.72% 79.01%\93.03% R1-Distill-Verifier-1.5B 76.18%\81.22% 80.71%\86.30% 77.91%\78.57% 88.77%\89.50% 81.91%\86.36% General LLM as Judge Qwen2.5-7B-Instruct 77.88%\47.51% 89.45%\93.59% 54.62%\29.46% 86.66%\86.64% 82.35%\75.90% Qwen2.5-14B-Instruct 78.08%\59.12% 90.15%\86.01% 62.25%\34.82% 91.67%\92.47% 84.75%\80.21% Qwen2.5-32B-Instruct 81.08%\71.27% 85.28%\96.50% 62.25%\48.21% 95.39%\97.37% 81.20%\88.36% Qwen3-8B 70.77%\74.31% 88.04%\93.88% 81.53%\83.04% 95.09%\97.37% 84.38%\90.79% Qwen3-14B 85.39%\80.11% 92.61%\95.92% 84.34%\92.86% 96.99%\98.52% 91.11%\93.68% Qwen3-32B 74.67%\70.72% 83.82%\93.00% 83.40%\91.96% 95.89%\97.15% 84.69%\90.31% DS-R1-Distill-Qwen-7B 64.66%\74.59% 75.38%\85.42% 74.80%\77.68% 91.78%\97.03% 77.07%\88.72% DS-R1-Distill-Qwen-14B 76.18%\78.73% 81.06%\85.13% 68.67%\64.29% 95.69%\97.26% 83.02%\88.66% DS-R1-Distill-Qwen-32B 72.37%\78.73% 79.10%\96.50% 72.29%\85.71% 96.59%\99.43% 81.85%\93.50%
Boxed-only Input and Short Output.
[…]
Full-CoT Input and Short Output.
Here, the verifier receives the full reasoning process as input but remains limited in output length. This setting challenges the model to efficiently extract conclusions from complex context. Specialized verifiers again show dominant performance, particularly in accuracy. For instance, xVerify-9B-C yields the best overall accuracy (90.96%) and leads in most subjects. While general LLMs such as Qwen2.5-7B-Instruct demonstrate high recall (e.g., 91.99% in math), they often trade precision for inclusiveness. This highlights a key contrast: general LLMs can tolerate diverse answer forms but struggle with precision. The details are shown in Table 2.
Boxed-only Input and Long Output.
This setting introduces detailed verifier responses while maintaining a simplified input. It evaluates the reasoning and explanatory abilities under minimal input. Figure 5 displays the detailed results. Specialized verifiers like xVerify-9B-C and R1-Distill-Verifier-1.5B continue to lead in accuracy and show robust recall, with the latter achieving 93.98% overall recall. While general models (e.g., Qwen3-14B) exhibit impressive recall in physics and chemistry, their accuracy often falls short due to misjudgment of semantically inconsistent responses.
Full-CoT Input and Long Output.
This most comprehensive setting approximates real-world judgment scenarios, allowing verifiers to access both rich context and full expressive capacity. Details are presented in Table 3. Among all models, xVerify-9B-C again leads with top-tier accuracy across chemistry (96.18%), biology (89.96%), and physics (93.98%). Its strong performance underscores the effectiveness of structured training for specialized verifiers. Meanwhile, Qwen3-14B, a general LLM, achieves comparable overall accuracy (91.11%) and excels in recall for subjects like physics and biology. Nonetheless, specialized verifiers maintain an advantage in precision and consistency.
Analysis
Performance Across Settings.
We explore two input types (extracted vs. complete) and two output lengths (8 vs. 4k tokens), reflecting various levels of evaluation granularity. Across all configurations, specialized verifiers consistently achieve higher accuracy, especially in fields demanding strict semantic consistency. Richer input and longer output lead to performance gains, but recall improvements are modest—indicating a preference for precision over inclusiveness. For instance, xVerify models maintain high accuracy in various settings, demonstrating strong reliability. In contrast, general LLMs exhibit greater sensitivity to input/output conditions. They perform poorly when limited to short outputs and simple inputs but show substantial gains when given more context and freedom. Larger LLMs (e.g., Qwen3-14B/32B, DeepSeek-R1-Distill-Qwen-32B) exhibit huge recall improvements, sometimes surpassing specialized models. However, their accuracy remains less stable due to looser judgment standards and overgeneralization.
Strictness vs. Inclusiveness.
Specialized verifiers prioritize correctness and reject ambiguous or loosely matched responses, aiming to reduce false positives. This yields high accuracy but may sacrifice recall, particularly when faced with valid answer variants. In contrast, general LLMs adopt a more inclusive stance, recognizing broader expression forms and redundant reasoning. While this boosts recall, it increases the risk of accepting incorrect answers.
Towards End-to-End Judgment.
From a system perspective, verifiers should ideally produce direct, structured outputs (e.g., “Correct”/“Incorrect”) without relying on intermediate answer extraction or post-processing. This reduces engineering overhead and minimizes error propagation. Specialized verifiers show promise in supporting this end-to-end evaluation paradigm. Their structured training enables them to focus on key conclusions within noisy or verbose outputs. Enhancing robustness to formatting and expression variance is a promising direction.
Recommendations and Hybrid Strategies.
Specialized models can benefit from training augmentation with varied answer forms and noisy reasoning chains to boost generalization and recall. General LLMs should be guided toward more structured outputs and fine-tuned for domain-specific judgment accuracy. A hybrid pipeline is recommended: use general LLMs for high-recall coarse filtering, followed by specialized verifiers for precision filtering—balancing flexibility with accuracy.
Conclusion
This benchmark presents a comprehensive evaluation of specialized verifiers and general LLMs under varying input-output constraints across multiple subjects. Results show that specialized verifiers consistently lead in accuracy, particularly in structured tasks, while larger general LLMs excel in recall and flexibility, especially when given full input and output freedom. However, both approaches have trade-offs: specialized verifiers struggle with answer diversity, and general LLMs risk misjudgment due to less precise criteria and unstructured outputs. Moreover, answer format and extraction dependency impose challenges for practical deployment. To address these gaps, future models should aim for end-to-end judgment with minimal reliance on answer extraction, while improving robustness to expression variability. We recommend enhancing generalization in specialized verifiers and guiding general LLMs toward structured outputs through targeted fine-tuning.

### v3

@@
 To address this, we propose VerifyBench–a cross-domain comprehensive benchmark for systematically evaluating verifiers.
-We construct 4,000 expert-level questions covering mathematics, physics, chemistry, and biology.
+We construct about 4,000 expert-level questions covering mathematics, physics, chemistry, and biology.
 Each question is equipped with reference answers and diverse responses.
@@
 To fill this critical void, we introduce VerifyBench: a cross-domain comprehensive benchmark for systematic evaluation of verifiers.
-VerifyBench is constructed from 4,000 expert-level questions spanning mathematics, physics, chemistry, and biology.
+VerifyBench is constructed from about 4,000 expert-level questions spanning mathematics, physics, chemistry, and biology.
 Each question includes reliable reference answers and diverse Chain-of-Thought (CoT) (Ma et al.
@@
 Our main contributions are as follows:
-We introduce VerifyBench, a novel, multidisciplinary benchmark of 4,000 expert-level questions across mathematics, physics, chemistry, and biology, featuring fine-grained human annotations and diverse CoT responses from state-of-the-art LLM, specifically designed for systematic verifier evaluation.
+We introduce VerifyBench, a novel, multidisciplinary benchmark of about 4,000 expert-level questions across mathematics, physics, chemistry, and biology, featuring fine-grained human annotations and diverse CoT responses from state-of-the-art LLM, specifically designed for systematic verifier evaluation.
 We conduct a comprehensive empirical study comparing various verifier types—specialized models vs.
@@
 Question Collection and Solution Generation
-We curate a total of 4,000 high-quality, expert-level questions across four disciplines: mathematics, physics, chemistry, and biology (1,000 per subject).
+We have carefully selected a total of approximately 4,000 high-quality, expert-level questions, covering four disciplines: mathematics, physics, chemistry, and biology, with around 1,000 questions per subject.
 These questions are designed to be challenging, incorporating rich domain-specific background, requiring advanced reasoning capabilities from LLMs, and allowing for open-ended, expressive answer spaces—thereby increasing benchmark complexity.
@@
 These questions are chosen to ensure broad topical coverage, high cognitive demand, and diversity in conceptual, theoretical, and applied reasoning, always paired with correct and clear reference answers.
-A detailed description of our question selection criteria and source identification process is provided in Appendix.
+A detailed description of our question selection criteria and source identification process is provided in the Appendix.
 Each question undergoes 2 rounds of manual review.
@@
 This careful review process ensures that the questions and reference answers are professional, coherent, and free from noise.
-Further specifics on our review protocols and reviewer qualifications can be found in Appendix.
+Further specifics on our review protocols and reviewer qualifications can be found in the Appendix.
 […]
@@
 Technical Formatting and Language:
-Tolerances are defined for spelling, punctuation, and formatting. In long-chain reasoning, correctness of logic and final answer are prioritized over surface-level language differences.
+Differences in spelling, punctuation, and formatting are acceptable. In long-chain reasoning, correctness of logic and final answer are prioritized over surface-level language differences.
 Benchmark Statistics
 This section offers a detailed summary of VerifyBench, highlighting its scale, composition, and core characteristics across disciplines.
-The benchmark contains 4,000 questions, each paired with a reference answer and a model-generated response, intended to support comprehensive and rigorous verifier evaluation.
-To ensure balanced representation, we include 1,000 questions from each of four major scientific domains: mathematics, physics, chemistry, and biology.
+The benchmark contains about 4,000 questions, each paired with a reference answer and a model-generated response, intended to support comprehensive and rigorous verifier evaluation.
+To ensure balanced representation, we include about 1,000 questions from each of four major scientific domains: mathematics, physics, chemistry, and biology.
 Key statistics are summarized in Table 1.
 Statistic | Value |
-Total Questions | 4,000 |
-Questions per Discipline | 1,000 |
+Total Questions | 3,989 |
 Average Question Length | 186 tokens |
 Average Model Response Length | 4,553 tokens |
-Total Annotated Instances | 4,000 |
+Total Annotated Instances | 3,989 |
 Label Distribution (Correct / Incorrect) | 45% / 55% |
 Inter-Annotator Agreement (IAA) | 0.88 – 0.92 |
-Verifier Math Chemistry Biology Physics Overall Trained Verifier xVerify-0.5B-I 79.60%\64.92% 94.77%\96.50% 86.75%\92.86% 92.48%\94.86% 88.77%\88.66% xVerify-3B-Ia 81.28%\56.91% 94.77%\96.21% 89.56%\90.18% 91.57%\92.24% 89.24%\85.35% xVerify-8B-I 81.58%\55.80% 96.39%\96.50% 87.95%\91.07% 92.68%\92.69% 90.03%\85.47% xVerify-9B-C 83.03%\54.14% 96.28%\94.46% 89.96%\91.07% 93.78%\92.81% 90.96%\84.76% general-verify - - - - - R1-Distill-Verifier-1.5B - - - - - General LLM as Judge Qwen2.5-7B-Instruct 78.18%\91.99% 89.45%\82.22% 68.67%\83.04% 86.66%\98.17% 83.52%\93.74% Qwen2.5-14B-Instruct 78.40%\90.61% 90.05%\82.80% 54.62%\83.04% 91.28%\96.46% 84.10%\91.55% Qwen2.5-32B-Instruct 80.98%\76.24% 85.63%\61.81% 61.45%\66.96% 95.39%\97.26% 85.34%\83.58% Qwen3-8B/14B/32B - - - - - DS-R1-Distill-Qwen-7B/14B/32B - - - - -
+Verifier Mathematics Chemistry Biology Physics Overall Trained Verifier xVerify-0.5B-I 79.60%\64.92% 94.77%\96.50% 86.75%\92.86% 92.48%\94.86% 88.77%\88.66% xVerify-3B-Ia 81.28%\56.91% 94.77%\96.21% 89.56%\90.18% 91.57%\92.24% 89.24%\85.35% xVerify-8B-I 81.58%\55.80% 96.39%\96.50% 87.95%\91.07% 92.68%\92.69% 90.03%\85.47% xVerify-9B-C 83.03%\54.14% 96.28%\94.46% 89.96%\91.07% 93.78%\92.81% 90.96%\84.76% general-verify - - - - - R1-Distill-Verifier-1.5B - - - - - General LLM as Judge Qwen2.5-7B-Instruct 78.18%\91.99% 89.45%\82.22% 68.67%\83.04% 86.66%\98.17% 83.52%\93.74% Qwen2.5-14B-Instruct 78.40%\90.61% 90.05%\82.80% 54.62%\83.04% 91.28%\96.46% 84.10%\91.55% Qwen2.5-32B-Instruct 80.98%\76.24% 85.63%\61.81% 61.45%\66.96% 95.39%\97.26% 85.34%\83.58% Qwen3-8B/14B/32B - - - - - DS-R1-Distill-Qwen-7B/14B/32B - - - - -
 VerifyBench emphasizes long-form, multi-step reasoning, representing a substantial challenge for existing verifiers.
-The high average token count for both responses (Table 1) reflects the benchmark’s focus on detailed, explanatory solutions and verbose model outputs.
+The high average token count for responses (Table 1) reflects the benchmark’s focus on detailed, explanatory solutions and verbose model outputs.
 Our two-stage human annotation pipeline ensures high-quality binary labels, with inter-annotator agreement (IAA) (Artstein 2017) scores ranging from 0.88 to 0.92, confirming annotation reliability.
@@
 Response Generation.
-To simulate realistic response patterns of reasoning-capable language models, we employ QwQ-32B (Team 2025) to generate detailed CoT responses for all 4,000 multi-subject questions in VerifyBench.
+To simulate realistic response patterns of reasoning-capable language models, we employ QwQ-32B (Team 2025) to generate detailed CoT responses for all multi-subject questions in VerifyBench.
 We use the official recommended hyperparameters: temperature 0.6, TopP 0.95, MinP 0, TopK 40, and no repetition penalty.
@@
 Baseline.
-We evaluate two broad categories of LLMs as verifiers.
+We evaluate two categories of LLMs as verifiers.
 The first category comprises general-purpose open-source LLMs, including the Qwen2.5-Instruct series (7B, 14B, 32B) (Qwen et al.
@@
 2025).
-The second category consists of specialized verifiers that have been trained on large-scale verification datasets, such as the xVerify series (0.5B, 3B, 8B, 9B) (Chen et al.
+The second category consists of specialized verifiers trained on large-scale verification datasets, such as the xVerify series (0.5B, 3B, 8B, 9B) (Chen et al.
 2025), general_verifier(Ma et al.
@@
 Among these, models from the xVerify and Qwen2.5-Instruct series directly output binary verification decisions (e.g., “correct” or “incorrect”).
-In contrast, generative models—such as the Qwen3 and DeepSeek-R1-Distill-Qwen series, as well as general_verifier and R1-Distill-Verifier-1.5B—produce reasoning traces alongside their final judgments.
+In contrast, generative models—such as the Qwen3 and DeepSeek-R1-Distill-Qwen series, as well as general_verifier and R1-Distill-Verifier-1.5B—produce reasoning traces alongside final judgments.
 Verifier Inference.
@@
 Evaluation Variants.
-We consider the following four experimental settings:
+We consider the following four experimental settings, all of which include the question and reference answer as inputs:
 […]
@@
 Main Results
-Verifier Math Chemistry Biology Physics Overall Trained Verifier xVerify-0.5B-I 79.38%\54.14% 94.67%\94.46% 87.15%\91.07% 92.38%\92.81% 88.64%\84.76% xVerify-3B-Ia 81.18%\56.91% 95.03%\96.21% 88.76%\90.18% 91.88%\92.24% 89.38%\85.35% xVerify-8B-I 81.58%\55.80% 96.28%\96.50% 89.16%\91.07% 92.58%\92.69% 90.06%\85.47% xVerify-9B-C 82.78%\64.92% 96.18%\96.50% 89.96%\92.86% 93.98%\94.86% 90.86%\88.66% general-verify 68.77%\88.12% 75.13%\88.63% 73.09%\85.71% 94.32%\97.72% 79.01%\93.03% R1-Distill-Verifier-1.5B 76.18%\81.22% 80.71%\86.30% 77.91%\78.57% 88.77%\89.50% 81.91%\86.36% General LLM as Judge Qwen2.5-7B-Instruct 77.88%\47.51% 89.45%\93.59% 54.62%\29.46% 86.66%\86.64% 82.35%\75.90% Qwen2.5-14B-Instruct 78.08%\59.12% 90.15%\86.01% 62.25%\34.82% 91.67%\92.47% 84.75%\80.21% Qwen2.5-32B-Instruct 81.08%\71.27% 85.28%\96.50% 62.25%\48.21% 95.39%\97.37% 81.20%\88.36% Qwen3-8B 70.77%\74.31% 88.04%\93.88% 81.53%\83.04% 95.09%\97.37% 84.38%\90.79% Qwen3-14B 85.39%\80.11% 92.61%\95.92% 84.34%\92.86% 96.99%\98.52% 91.11%\93.68% Qwen3-32B 74.67%\70.72% 83.82%\93.00% 83.40%\91.96% 95.89%\97.15% 84.69%\90.31% DS-R1-Distill-Qwen-7B 64.66%\74.59% 75.38%\85.42% 74.80%\77.68% 91.78%\97.03% 77.07%\88.72% DS-R1-Distill-Qwen-14B 76.18%\78.73% 81.06%\85.13% 68.67%\64.29% 95.69%\97.26% 83.02%\88.66% DS-R1-Distill-Qwen-32B 72.37%\78.73% 79.10%\96.50% 72.29%\85.71% 96.59%\99.43% 81.85%\93.50%
+Verifier Mathematics Chemistry Biology Physics Overall Trained Verifier xVerify-0.5B-I 79.38%\54.14% 94.67%\94.46% 87.15%\91.07% 92.38%\92.81% 88.64%\84.76% xVerify-3B-Ia 81.18%\56.91% 95.03%\96.21% 88.76%\90.18% 91.88%\92.24% 89.38%\85.35% xVerify-8B-I 81.58%\55.80% 96.28%\96.50% 89.16%\91.07% 92.58%\92.69% 90.06%\85.47% xVerify-9B-C 82.78%\64.92% 96.18%\96.50% 89.96%\92.86% 93.98%\94.86% 90.86%\88.66% general-verify 68.77%\88.12% 75.13%\88.63% 73.09%\85.71% 94.32%\97.72% 79.01%\93.03% R1-Distill-Verifier-1.5B 76.18%\81.22% 80.71%\86.30% 77.91%\78.57% 88.77%\89.50% 81.91%\86.36% General LLM as Judge Qwen2.5-7B-Instruct 77.88%\47.51% 89.45%\93.59% 54.62%\29.46% 86.66%\86.64% 82.35%\75.90% Qwen2.5-14B-Instruct 78.08%\59.12% 90.15%\86.01% 62.25%\34.82% 91.67%\92.47% 84.75%\80.21% Qwen2.5-32B-Instruct 81.08%\71.27% 85.28%\96.50% 62.25%\48.21% 95.39%\97.37% 81.20%\88.36% Qwen3-8B 70.77%\74.31% 88.04%\93.88% 81.53%\83.04% 95.09%\97.37% 84.38%\90.79% Qwen3-14B 85.39%\80.11% 92.61%\95.92% 84.34%\92.86% 96.99%\98.52% 91.11%\93.68% Qwen3-32B 74.67%\70.72% 83.82%\93.00% 83.40%\91.96% 95.89%\97.15% 84.69%\90.31% DS-R1-Distill-Qwen-7B 64.66%\74.59% 75.38%\85.42% 74.80%\77.68% 91.78%\97.03% 77.07%\88.72% DS-R1-Distill-Qwen-14B 76.18%\78.73% 81.06%\85.13% 68.67%\64.29% 95.69%\97.26% 83.02%\88.66% DS-R1-Distill-Qwen-32B 72.37%\78.73% 79.10%\96.50% 72.29%\85.71% 96.59%\99.43% 81.85%\93.50%
 Boxed-only Input and Short Output.
@@
 Full-CoT Input and Short Output.
-Here, the verifier receives the full reasoning process as input but remains limited in output length.
+The verifier receives the full reasoning process as input but remains limited in output length.
 This setting challenges the model to efficiently extract conclusions from complex context.
@@
 Towards End-to-End Judgment.
-From a system perspective, verifiers should ideally produce direct, structured outputs (e.g., “Correct”/“Incorrect”) without relying on intermediate answer extraction or post-processing.
+From a system perspective, verifiers should ideally produce direct, structured outputs (e.g., “Correct”/“Incorrect”) without relying on the extraction of model response or verifier’s judgment result.
 This reduces engineering overhead and minimizes error propagation.
@@
 Recommendations and Hybrid Strategies.
-Specialized models can benefit from training augmentation with varied answer forms and noisy reasoning chains to boost generalization and recall.
+Specialized verifiers can benefit from training augmentation with varied answer forms and noisy reasoning chains to boost generalization and recall.
 General LLMs should be guided toward more structured outputs and fine-tuned for domain-specific judgment accuracy.
-A hybrid pipeline is recommended: use general LLMs for high-recall coarse filtering, followed by specialized verifiers for precision filtering—balancing flexibility with accuracy.
+A hybrid pipeline is recommended: use general LLMs for high-recall coarse filtering, followed by specialized verifiers for precision filtering.
 Conclusion
 This benchmark presents a comprehensive evaluation of specialized verifiers and general LLMs under varying input-output constraints across multiple subjects.
-Results show that specialized verifiers consistently lead in accuracy, particularly in structured tasks, while larger general LLMs excel in recall and flexibility, especially when given full input and output freedom.
+Results show that specialized verifiers consistently lead in accuracy, particularly in structured tasks, while general LLMs excel in recall and flexibility, especially when they have a larger model size and are given full input and output freedom.
 However, both approaches have trade-offs: specialized verifiers struggle with answer diversity, and general LLMs risk misjudgment due to less precise criteria and unstructured outputs.
 Moreover, answer format and extraction dependency impose challenges for practical deployment.
-To address these gaps, future models should aim for end-to-end judgment with minimal reliance on answer extraction, while improving robustness to expression variability.
-We recommend enhancing generalization in specialized verifiers and guiding general LLMs toward structured outputs through targeted fine-tuning.
+To address these gaps, future models should aim for end-to-end judgment with minimal or no reliance on the extraction of model-generated answers and verifiers’ judgment results, while improving robustness to expression variability.
+We recommend enhancing generalization in specialized verifiers and guiding general LLMs toward structured outputs through fine-tuning.